OpenStudy (anonymous):

I just need the answer, no explanation. use part 1 of the fundamental thm of calc to find the derivative f(x)=integral from -5 to x of sqrt(t^3+125)dt

hartnn (hartnn):

why do u need only answers? have u solved it and verifying ?

hartnn (hartnn):

because here its against code of conduct to just give answers.

OpenStudy (anonymous):

Ans: sqrt(t^3+125)

hartnn (hartnn):

for this you can use \(\huge \frac{d}{dx}\int \limits_a^x f(t)dt=f(x)\)

OpenStudy (anonymous):

I got sqrt(x^3+125) is that right?

hartnn (hartnn):

yes. thats f(x).

hartnn (hartnn):

means its correct.